OpenStudy (anonymous):

HELP I GIVE MEDAL The area of a poster board is x2 + 3x - 10 inches. The width is x - 2 inches. Select the polynomial for the area of the poster board if one inch is removed from each side.

OpenStudy (anonymous):

Formula : Area=length * width Factor x^2+3x-10=(x+5)(x-2) then Length=(x+5) if one inch is removed from each side then the width becomes (x-3) and the length becomes (x+4) then using FOIL you will get. (x-3)(x+4)= x^2 +x -12 so it's x^2 + x - 12

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

welcome, anytime.